# Medium effects on the electromagnetic form factors of $\rho$ meson

### v1

@@ VII Summary and Conclusion
As a result, we have found that the $\rho$ meson effective mass decreases as the nuclear matter density increases. This is consistent with that obtained in Ref. [52], although the magnitude is different. We find the $\rho$ meson mass reduction at $\rho_{B}/\rho_{0}=1.0$ relative to that in free space is about 10%, which is consistent with that obtained in the QCD sun rule (QCDSR) of Ref. [43], and a very recent analysis in the inverse QCD sum rule (IQCDSR) of Ref. [45], giving about (10-20)%. However, the reduction of the $m_{\rho}^{*}$ of Ref. [52] is rather different from that obtained in the present work, as clearly shown in Fig. 2(a), where in the former, the in-medium $\rho$ mass was calculated in the quark-meson coupling (QMC) model using a rather larger free space light quark mass value compared to that used in the standard QMC model.
A similar trend is found in the in-medium dynamical quark mass, where the in-medium dynamical quark mass decreases as the nuclear matter density increases, as clearly illustrated in Fig. 2(b). The in-medium constituent quark mass results reported in Ref. [52] exhibit a similar trend to our predictions, showing a decrease with increasing nuclear matter density. However, their results show a more rapid decrease in the in-medium constituent quark mass calculated by the QMC model with a large free space light quark mass value.
@@
In addition, we have predicted the in-medium $\rho^{+}$ meson charge radius, that increases as the nuclear matter density increases. Our prediction for the in-medium charge radius is consistent with that obtained in Ref. [52], although different in magnitude, as indicated in Fig. 6(d).
Finally, we have predicted the $\rho^{+}$ meson magnetic and quadrupole moments in symmetric nuclear matter. We have found that the $\mu_{\rho}^{*}$ decreases as the nuclear matter density increases. The trends of our results are consistent with those obtained in Ref. [52], but different in magnitude. For the quadrupole moment, we have found that the $\mathcal{Q}_{\rho}^{*}$ increases as the nuclear matter density increases. Our predictions are consistent with those obtained in Ref. [52], however, again, different in magnitude, in particular at higher nuclear matter densities.
Overall, the predictions presented in this work are potentially testable by the future experiments, such as JPARC E16 [60], and CLAS at JLAB [39, 40], as well as lattice QCD [61, 62, 63]. For future work, we plan to apply the present approach to investigate the electromagnetic form factors of other spin-1 mesons in a nuclear medium, as well as pseudoscalar $D$ and $B$ mesons with heavy $c$ and $b$ quark contents, where such studies may require significant challenges, since heavy quark symmetry differs fundamentally from the light quark symmetry. Furthermore, we hope that the present study can provide useful guidance for lattice QCD in computing the electromagnetic form factors of the $\rho$ meson in free space as well as in a nuclear medium.
Acknowledgements
This work was supported by the PUTI Q1 Research Grant from the University of Indonesia (UI) under contract No. NKB 442/UN2.RST/HKP.05.00/2024. K.T. was supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq, Brazil), Processes No. 304199/2022-2, and FAPESP Process No. 2023/07313-6, and his work was also part of the projects, Instituto Nacional de Ciência e Tecnologia - Nuclear Physics and Applications (INCT-FNA), Brazil, Process No. 464898/2014-5.

### v2

@@ VII Summary and Conclusion
As a result, we have found that the $\rho$ meson effective mass decreases as the nuclear matter density increases. This is consistent with that obtained in Ref. [deMelo:2018hfw], although the magnitude is different. We find the $\rho$ meson mass reduction at $\rho_{B}/\rho_{0}=1.0$ relative to that in free space is about 10%, which is consistent with that obtained in the QCD sun rule (QCDSR) of Ref. [Hatsuda:1991ez], and a very recent analysis in the inverse QCD sum rule (IQCDSR) of Ref. [Mutuk:2025lak], giving about (10–20)%. However, the reduction of the $m_{\rho}^{*}$ of Ref. [deMelo:2018hfw] is rather different from that obtained in the present work, as clearly shown in Fig. 2(a), where in the former, the in-medium $\rho$ mass was calculated in the quark-meson coupling (QMC) model using a rather larger free space light quark mass value compared to that used in the standard QMC model.
A similar trend is found in the in-medium dynamical quark mass, where the in-medium dynamical quark mass decreases as the nuclear matter density increases, as clearly illustrated in Fig. 2(b). The in-medium constituent quark mass results reported in Ref. [deMelo:2018hfw] exhibit a similar trend to our predictions, showing a decrease with increasing nuclear matter density. However, their results show a more rapid decrease in the in-medium constituent quark mass calculated by the QMC model with a large free space light quark mass value.
@@
In addition, we have predicted the in-medium $\rho^{+}$ meson charge radius, that increases as the nuclear matter density increases. Our prediction for the in-medium charge radius is consistent with that obtained in Ref. [deMelo:2018hfw], although different in magnitude, as indicated in Fig. 6(d).
Finally, we have predicted the $\rho^{+}$ meson magnetic and quadrupole moments in symmetric nuclear matter. We have found that the $\mu_{\rho}^{*}$ decreases as the nuclear matter density increases. The trends of our results are consistent with those obtained in Ref. [deMelo:2018hfw], but different in magnitude. For the quadrupole moment, we have found that the $\mathcal{Q}_{\rho}^{*}$ increases as the nuclear matter density increases. Our predictions are consistent with those obtained in Ref. [deMelo:2018hfw], however, again, different in magnitude, in particular at higher nuclear matter densities.
Overall, the predictions presented in this work are potentially testable by the future experiments, such as JPARC E16 [Aoki:2015qla], and CLAS at JLAB [CLAS:2008jqp, CLAS:2007dll], as well as lattice QCD [Hedditch:2007ex, Owen:2015gva, Wang:2025hew]. For future work, we plan to apply the present approach to investigate the electromagnetic form factors of other spin-1 mesons in a nuclear medium, as well as pseudoscalar $D$ and $B$ mesons with heavy $c$ and $b$ quark contents, where such studies may require significant challenges, since heavy quark symmetry differs fundamentally from the light quark symmetry. Furthermore, we hope that the present study can provide useful guidance for lattice QCD in computing the electromagnetic form factors of the $\rho$ meson in free space as well as in a nuclear medium.
Acknowledgments
This work was supported by the PUTI Q1 Research Grant from the University of Indonesia (UI) under contract No. NKB 442/UN2.RST/HKP.05.00/2024. K.T. was supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq, Brazil), Processes No. 304199/2022-2, and FAPESP Process No. 2023/07313-6, and his work was also part of the projects, Instituto Nacional de Ciência e Tecnologia–Nuclear Physics and Applications (INCT-FNA), Brazil, Process No. 464898/2014-5.
Data availability
The data are not publicly available. The data are available from the authors upon reasonable request.